Mole fraction of solute in benzene is 0.2. Calcualte its molality of solute.
need the answer very urgentlyy!!!

Grade:11

3 Answers

Ashutosh Mohan Sharma
8 years ago
mole fraction = 0.2.
therefore in 100 moles, 20 moles are of solute and 80 moles are of benzene(solvent)
mass of benzene = 80 * Mol.Wt of Benzene = 80* 78 = 6240g
for 6240 g of solvent we have 20 moles of solute
Therefore for 1000g of solvent we have (20)*(1000)/(6240) = 3.20 molal
Aparna
6 years ago
no. of moles of solute/no. of moles of benzene =0.2=02/10=1/5.It means 1 mole solute in 4 moles benzeneMolality=no. of moles of solute*1000/weight of benzene in g.....(1).weight of benzene= no. of moles* molar mass=4*78=312Molality =1*1000/312=3.2 molal
Kushagra Madhukar
3 years ago
Dear student,
Please find the solution to your problem below.
 
mole fraction = 0.2.
therefore in 100 moles, 20 moles are of solute and 80 moles are of benzene(solvent)
mass of benzene = 80 * Mol.Wt of Benzene = 80* 78 = 6240g
for 6240 g of solvent we have 20 moles of solute
Therefore for 1000g of solvent we have (20)*(1000)/(6240) = 3.20 molal
 
Hope it helps.
Thanks for the question.
Regards,
Kushagra